OpenStudy (anonymous):

JUST NEEDED TO BE CHECKED! Solve the system x + y = 6 and x - y = 2. What is the value of x? OPTIONS: 6 2 4 0 WILL GET A MEDAL!

OpenStudy (anonymous):

just need it to be checked!

jimthompson5910 (jim_thompson5910):

What did you get

OpenStudy (anonymous):

I got ..

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

that's not correct, this might help x+y = 6 x-y = 2 Add the equations to get x+y = 6 x-y = 2 --------- 2x+0y = 8 So 2x = 8 which means x = ???

OpenStudy (anonymous):

Its four! dang it... the way I did was plug 2 in for X, and then guessed y... so... 2 + 4 (my guessed y) = 6 which worked 2-4 (my guessed y) = 2... I now just realized that it was negative 2... i did divided by... ohh. stupid me... sorry

OpenStudy (anonymous):

Please dont tell me i got it wrong again! :(

jimthompson5910 (jim_thompson5910):

yeah that gets a lot of people 2-4 = -2 not 2-4 = 2 technically they both have the same magnitude, just different sign just keep 2-4 in mind to mean "you have $2 and you spend $4 to have the balance land to -2 dollars...ie you are 2 dollars in debt"

jimthompson5910 (jim_thompson5910):

no you are correct, x = 4, y = 2 x+y = 6 4+2 = 6 6=6... works x-y = 2 4-2=2 2=2...works

OpenStudy (anonymous):

I feel stupid... thanks so much! if I could give you 1000 medals i would... but I cant I became a fan though! So thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

one medal is fine, but I appreciate the thought